RC meeting and Run Plan, Thursday November 4th

From SBS wiki
Jump to: navigation, search

Daily_Meetings

RC Meeting Information

RC Daily Meetings are at 3:15pm, remote connection link: https://bluejeans.com/256205934

Accelerator Status

The accelerator is still intending to be able to send beam to Hall A by 8:30 am on Monday if we are able to receive it.

Hall A Status and News

Jessie still does not have signed paperwork while will allow the movement of SBB. It is therefore very unlikely that Hall A will be ready to take beam on Monday.


Subsystem Updates

  1. HCAL: Verify led signals. Will tears new pattern. Efficiency. Implement tdc alignment parameters.
  2. BBCal: Ellen ready with sensors. 9 on hand. 2 on each side of spectrometer. Will monitor in epics. Cosmic overnight.
  3. Timing horoscope: refining the calibration. Preliminary timewalk correction improved resolution by 100 ps.
  4. Grinch: led run overnight, looks normal now. HV probably too high on some channels. Pulse rate increased to 100hz.
  5. GEMS: back on N2 to conserve gas. Sean analyzed comic data. Zeke will fix LV in hall.
  6. Beamline: trouble with reply, probably clock problem. SBS side seems okay.
  7. DAQ: seems okay. Run all detectors tonight.
  8. Replay Analysis: update to GEM code to shrink memory footprint. Minor changes to offline zero suppression and software threshold in cluster formation. Merged in. Working to be ready for first replay. Optics analysis ongoing. Calibration and resolution studios ongoing. GEM occupancy studies ongoing.